### v3

# Chains without regularity

Abstract
We study chain-recurrence and chain-transitivity in compact dynamical systems without any regularity assumptions. We prove that every compact system has a chain-recurrent point and a closed, invariant, chain-transitive subsystem. We also prove that, in fact, the same results hold for a strengthened form of chain-recurrence, which we identify.
KEYWORDS: Dynamical systems; Topological dynamics; Chain-recurrence; Discontinuous dynamics.
MSC2020: 37B20, 37B65.
1 Introduction
Chain-recurrence is arguably the most important concept in topological dynamics. On one hand, it has strong connections with applications. Any numerical/experimental evaluation of the time-evolution of a system has nonzero round-off/observational errors bounded by some $\epsilon>0$, so what it really provides is an $\epsilon$-chain rather than an orbit; a theoretical understanding of $\epsilon$-chains is thus crucial for linking abstract dynamical results to what can be in fact computed or measured. On the other hand, the concept of chain-recurrence has deep consequences for our theoretical understanding of long-term behaviors of dynamical systems, underpinning in particular C. Conley’s decomposition results ([1]) and E. Akin’s characterization of attractors ([2], p. 82).
The aim of this work is to show that chain-recurrence and chain-transitivity retain considerable significance even without assuming that the evolution map is continuous. In fact, key existence theorems can be obtained, not merely when relaxing continuity to weakened forms of regularity (such as quasi-continuity, Baire class 1 or a.e. continuity), but also for utterly arbitrary maps. Moreover, the proofs are constructive, in the sense that they do not require any form of the Axiom of Choice. In broader terms, this can be summarized by saying that the constraints imposed by compactness and chain relations prevail over the potentially unbounded set-theoretic complexity implied by arbitrary maps.
Let us start by recalling a well-known fact: consider a dynamical system $(X,f)$, with $X$ a metrizable space. When $X$ is compact and $f$ is continuous, a classical result by Birkhoff ([3]) ensures that there is at least one $x\in X$ that revisits each of its neighborhoods within a bounded number of iterations, that is there exists an almost-periodic point. If we drop the continuity assumption, this is obviously false in general. However, one can wonder if a point verifying some more general form of self-recurrence always exists even without continuity. Searching for a suitable candidate generalization of almost-periodicity, let us review the well-known topological dynamical relations $\mathcal{O},\mathcal{R},\mathcal{N},\mathcal{C}\subseteq X^{2}$, as introduced by E. Akin ([4]).
For $d:X^{2}\to\mathbb{R}_{0}^{+}$ a metric compatible with the topology on $X$, the topological dynamical relations can be defined as follows (see for instance [2, Def. 2.2, p. 47]):
Definition 1.1.
  1. 1.
    Orbit relation:
    $x\,\mathcal{O}\,y$ iff $\exists\,k\in\mathbb{N}\,|\,f^{k}(x)=y$;
  2. 2.
    Recurrence relation:
    $x\,\mathcal{R}\,y$ iff $\forall\epsilon>0\,\exists\,k\in\mathbb{N}\,|\,f^{k}(x)\in B_{\epsilon}(y)$;
  3. 3.
    Non-wandering relation:
    $x\,\mathcal{N}\,y$ iff $\forall\epsilon>0\,\exists\,z\in B_{\epsilon}(x)\,\exists\,k\in\mathbb{N}\,|\,f^{k}(z)\in B_{\epsilon}(y)$;
  4. 4.
    Chain-recurrence relation:
    $x\,\mathcal{C}\,y$ iff $\forall\epsilon>0\,\exists\,n\in\mathbb{N},\,\exists\,x_{0},\dots,x_{n}\,|x_{0}=x,x_{n}=y,\,d(f(x_{i}),x_{i+1})<\epsilon\ \ (i=0,\dots,n-1)$.
Notice that all the relations, although defined in metric terms, are independent of the metric.
By compactness, every infinite orbit has an accumulation point, which is in fact non-wandering, in the sense that it verifies $x\,\mathcal{N}\,x$. Indeed, any open neighborhood of the accumulation point $x$ of an infinite orbit $\mathcal{O}(y)$ contains two points of type $f^{k}(y)$, $f^{h}(y)$ for some distinct natural numbers $h,k$, and thus the claim in Def.1.1-3 is verified with $y=x$. Since finite orbits contain periodic (hence non-wandering) points, we have the following, easy
Fact.
If $X$ is compact, the non-wandering set $\{x\in X|x\,\mathcal{N}\,x\}$ of $(X,f)$ is nonempty, independently of any assumption on $f$.
However, the significance of this result is problematic if we cannot guarantee that $f$ is continuous at the non-wandering point $x$. Specifically, the statement $x\,\mathcal{N}\,x$ is independent of the value of $f$ at $x$, so it provides insight into the dynamics of $f$ only when the behavior of $f$ at $x$ can be inferred from its values at sufficiently close points — that is, when $f$ is continuous.
A consequence of this is that, without regularity assumptions, in general we have $\mathcal{N}\not\subseteq\mathcal{C}$, so that, in particular, it does not directly follow from Fact that every compact system has a chain-recurrent point. Luckily, it is easy to modify slightly the definition of $\mathcal{N}$ introducing a relation $\widetilde{\mathcal{N}}$ such that:
  • $\widetilde{\mathcal{N}}=\mathcal{N}$ in case $f$ is continuous;
  • the points in $\widetilde{\mathcal{N}}$-relation with $x$ depend, in general, on the value of $f$ at $x$;
  • we have $\widetilde{\mathcal{N}}\subseteq\mathcal{C}$ also without continuity.
If $f$ is continuous, indeed, it does not matter whether we make the $\epsilon$-correction before or after the first iteration of the map, so in the continuous case the relation $\widetilde{\mathcal{N}}$ defined below is just equivalent to $\mathcal{N}$:
Definition 1.2.
For $x,y\in X$, we write $x\,\widetilde{\mathcal{N}}\,y$ if for every $\epsilon>0$ there exist $k\in\mathbb{N}_{0}$ and $z\in B_{\epsilon}(f(x))$ such that $f^{k}(z)\in B_{\epsilon}(y)$.
In case of a discontinuous map, Def.1.2 is not anymore equivalent to Def.1.1-3. In fact, the new version fits better with the other topological dynamical relations. In particular, the statement $x\,\mathcal{A}\,x$ depends on the behavior of $f$ at $x$ when $\mathcal{A}$ is any of the relations $\mathcal{O}$, $\mathcal{R}$, $\widetilde{\mathcal{N}}$, $\mathcal{C}$, and we have always $\mathcal{O}\subseteq\mathcal{R}\subseteq\widetilde{\mathcal{N}}\subseteq\mathcal{C}$ (this is graphically represented in the scheme of Fig.1), which resembles what happens in the continuous case, when $\mathcal{N}=\widetilde{\mathcal{N}}\subseteq\mathcal{C}$.
Figure 1: Topological dynamical relations. It is graphically emphasized that, if we replace $\mathcal{N}$ by $\widetilde{\mathcal{N}}$, additional $\epsilon$-corrections are allowed whenever one goes down in the table (right).
Let us now come back to our original problem: the existence of non-wandering points. If we do not assume continuity, it is false that a compact system has always a $\widetilde{\mathcal{N}}$-recurrent point. Indeed, an elementary counterexample to the claim that every compact metrizable dynamical system has a point $x$ such that $x\,\widetilde{\mathcal{N}}\,x$ is given by $([0,1],f)$ where $f$ is defined as:
$f(x)=\begin{cases}\frac{3}{4}&\quad\text{ if }x=0\\ \\ x(x+\frac{1}{2})&\quad\text{ if }x\in(0,\frac{1}{2})\\ \\ \frac{1}{4}&\quad\text{ if }x=\frac{1}{2}\\ \\ \frac{1}{2}x+\frac{1}{4}&\quad\text{ if }x\in(\frac{1}{2},1]\end{cases}$ (1)
Since all the topological relations $\mathcal{O},\mathcal{R},\widetilde{\mathcal{N}},\mathcal{C}$ are sensitive to what happens pointwise, one may think that there is little hope of finding any relevant recurrence phenomenon if the map is completely general. This impression is strengthened by the fact that, in the example above, $\widetilde{\mathcal{N}}$ is ruled out by an interval map with just two discontinuity points. If one allows arbitrary maps, the images of $\epsilon$-balls, and in general of open sets, can have arbitrary high (descriptive) set-theoretic complexity, so it seems a priori difficult that any topologically characterized recurrence phenomenon is forced to emerge.
Remarkably, this is not the case: the chain-recurrence relation $\mathcal{C}$ is robust enough, and the compactness hypothesis is sufficiently strict, to ensure key existence results without requiring assumptions on the map. Indeed we will show that:
  1. 1.
    every compact system has a chain-recurrent point;
  2. 2.
    every compact system has a chain transitive subsystem;
  3. 3.
    every point in a compact system is in chain relation with a chain-recurrent point;
  4. 4.
    every point in a compact system is in chain relation with a chain transitive subsystem.
Notice that points 2. and 4. are respectively stronger than 1. and 3., but their proof builds on their weaker versions, and thus for clarity we preferred to present them as separate, warm-up results. Finally, we will also show that, in fact, the chain-recurrent point whose existence is established is also recurrent with respect to a stronger relation than chain-recurrence, which we identify.
2 Settings
We say that $(X,f)$ is a compact dynamical system (or simply a dynamical system) if $X$ is a compact, metrizable space and $f$ is a self-map of $X$.
For $x\in X$, we indicate by $\mathcal{O}(x)$ the orbit of $x$, that is the set $\mathcal{O}(x)=\{f(x),f^{2}(x),\ldots\}$. We set $\mathcal{O}^{0}(x):=\{x\}\cup\mathcal{O}(x)$. We will indicate by $\mathbb{N}$ and $\mathbb{N}_{0}$, respectively, the set of positive and non-negative integers. Let us recall the notion of $\epsilon$-chain or $\epsilon$-pseudo-orbit ([2, p. 48]).
Definition 2.1.
Let $(X,f)$ be a dynamical system and $d:X^{2}\to\mathbb{R}_{0}^{+}$ be a metric compatible with the topology on $X$. Given two points $x,y\in X$ and $\epsilon>0$, an $\epsilon$-chain from $x$ to $y$ is a finite set of points $x_{0},x_{1},\ldots,x_{n}$ in $X$, with $n\geq 1$, such that
  1. i)
    $x_{0}=x$ and $x_{n}=y$,
  2. ii)
    $d(f(x_{i}),x_{i+1})<\epsilon$ for every $i=0,1,\ldots n-1$.
The chain relation $\mathcal{C}\subseteq X^{2}$, the binary relation mentioned in the Introduction, can be equivalently defined as follows ([2, Def. 2.2, p. 47]): given $x,y\in X$,
$x\,\mathcal{C}\,y\iff\forall\epsilon>0\text{ there exists an $\epsilon$-chain from $x$ to $y$}.$
Let $\mathcal{A}\subseteq X^{2}$ be a binary relation on $X$. For $M\subseteq X$ and $y\in X$, we write $M\,\mathcal{A}\,y$ if $x\,\mathcal{A}\,y$ for every $x\in M$. A point $x\in X$ is called chain-recurrent if $x\,\mathcal{C}\,x$. The set of all the chain-recurrent points of a dynamical system $(X,f)$ is denoted by $CR(X,f)$. We write simply $CR_{f}$ when the space $X$ is clear from the context.
Let us also recall the notion of strong chain-recurrence introduced by Easton [5].
Definition 2.2.
Let $(X,f)$ be a dynamical system and $d:X^{2}\to\mathbb{R}_{0}^{+}$ be a metric compatible with the topology on $X$. Given two points $x,y\in X$ and $\epsilon>0$, we say that a finite sequence of points $x_{0},x_{1},\ldots,x_{n}$ of $X$, with $n\in\mathbb{N}$, is a strong $(\epsilon,d)$-chain from $x$ to $y$ if
  1. i)
    $x_{0}=x$ and $x_{n}=y$,
  2. ii)
    $\sum_{i=0}^{n-1}d(f(x_{i}),x_{i+1})<\epsilon$.
The strong chain relation $\mathcal{SC}_{d}\subseteq X^{2}$ is the binary relation defined as follows: given $x,y\in X$,
$x\,\mathcal{SC}_{d}\,y\iff\forall\epsilon>0\text{ there exists a strong $(\epsilon,d)$-chain from $x$ to $y$}.$
A point $x\in X$ is called strong chain-recurrent if $x\,\mathcal{SC}_{d}\,x$. Let $\mathcal{SCR}_{d}(f)$ be the set of all the strong chain-recurrent points of $(X,f)$. We write simply $\mathcal{SCR}_{d}$ when the map $f$ is clear from the context.
The relation $\mathcal{SC}_{d}$ depends, in general, on the metric $d$ (while $\mathcal{C}$ can be defined in purely topological terms if the space is compact). Intersecting over all topological equivalent metrics on $X$, we obtain the generalized recurrent set $GR(f)$ introduced by Auslander for flows ([6]) and then extended to maps ([4, 7, 8]). More precisely, we give the following
Definition 2.3.
The generalized recurrent set, denoted by $GR(f)$, is defined as
$GR(f)=\bigcap_{d}\mathcal{SCR}_{d}(f),$
where the intersection is over all metrics $d$ compatible with the topology of $X$.
The following inclusions hold without any assumptions on $f$:
$GR(f)\subseteq\mathcal{SCR}_{d}(f)\subseteq CR_{f}.$
Definition 2.4.
Let $S$ be a subset of a topological space $X$. The derived set of $S$, denoted by $S^{\prime}$, is the set of the limit points $x$ of $S$, that is, the set of points $x$ such that for every neighborhood $U$ of $x$ we have that $S\cap(U\setminus\{x\})\neq\emptyset$.
Definition 2.5.
Let $(X,f)$ be a dynamical system, and $S\subseteq X$. We say that the set $S$ is $f$-invariant (or simply invariant) if $f(S)\subseteq S$.
Definition 2.6.
Let $(X,f)$ be a dynamical system. For $Y\subseteq X$, we say that $(Y,f|_{Y})$ is a subsystem of $(X,f)$ if $Y$ is closed and invariant.
Definition 2.7.
Let $(X,f)$ be a dynamical system, and $M\subseteq X$. We say that $M$ is chain-transitive ($\mathcal{C}$-transitive) if for every $x,y\in M$ we have that $x\,\mathcal{C}\,y$.
Definition 2.8.
Let $(X,f)$ be a dynamical system and $d:X^{2}\to\mathbb{R}_{0}^{+}$ be a metric compatible with the topology on $X$. We say that $M\subseteq X$ is strong chain-transitive ($\mathcal{SC}_{d}$-transitive) if for every $x,y\in M$ we have that $x\,\mathcal{SC}_{d}\,y$.
The following is a strengthening of chain-transitivity which will prove useful later on.
Definition 2.9.
Let $(X,f)$ be a dynamical system, and $M\subseteq X$. We say that $M$ is internally chain transitive (see e.g. [9, 10, 11], who use the concept in the case of continuous maps and flows) if for every $x,y\in M$ and for every $\epsilon>0$ there exists an $\epsilon$-chain form $x$ to $y$ whose points belong to $M$.
The concept of internal chain transitivity is related to what is called, in case of continuous maps, the restriction property for topological dynamical systems (see for instance [12], Prop. 2.7.21), which ensures that, setting $CR(X,f)=S$, we have
$CR(S,f|_{S})=S.$
[…]
3 Existence results
Theorem 3.1.
Let $(X,f)$ be a compact dynamical system. Then $CR_{f}\neq\emptyset$.
Proof.
If there is $x\in X$ such that $|\mathcal{O}(x)|<\aleph_{0}$, then at least one point $y$ in the orbit of $x$ is periodic, so that $y\,\mathcal{C}\,y$. We can therefore assume that every point in $X$ has an infinite orbit. Let $d:X^{2}\to\mathbb{R}_{0}^{+}$ be a metric compatible with the topology on $X$.
  • Step 1
    Pick $x\in X$ and set $x_{0}:=x$. For $k\in\mathbb{N}$, set $x_{k}:=f(x_{k-1})$. Pick $y\in\mathcal{O}(x)^{\prime}$ (which is nonempty by compactness) and set $x_{\omega}:=y$. Let us now define a set of points by transfinite induction as follows. Let $\alpha>\omega$ be an ordinal number. Assume that $x_{\beta}$ has been defined for every $\beta<\alpha$ and set:
    $\{x_{\eta}\}_{\beta\leq\eta<\alpha}:=S_{\beta,\alpha}(x)\quad,\quad S_{\alpha}(x):=S_{0,\alpha}(x).$
    1. 1.
      If $\alpha$ is a successor ordinal, set $x_{\alpha}:=f(x_{\alpha-1})$.
    2. 2.
      If $\alpha$ is a limit ordinal, then, recalling that every orbit in $X$ has infinite cardinality, we have $|S_{\beta,\alpha}(x)|\geq\aleph_{0}$ for every $\beta<\alpha$, so $\left(S_{\beta,\alpha}(x)\right)^{\prime}$ is closed and, by compactness, nonempty. Therefore,
      $S^{*}_{\alpha}(x):=\bigcap_{\beta<\alpha}\left(S_{\beta,\alpha}(x)\right)^{\prime}$ (2)
      is nonempty as well (see for instance [13, Th. 26.9, p. 169]). Pick then any $y\in S^{*}_{\alpha}(x)$ and set $x_{\alpha}:=y$.
  • Step 2
    For every ordinal $\alpha$, we have
    $x_{\beta}\,\mathcal{C}\,x_{\eta}\quad\text{ whenever }\quad 0\leq\beta<\eta\leq\alpha.$ (3)
    Let us prove this by induction on $\alpha$. If $\alpha=1$, the claim is obviously true. Assume then that $\alpha>1$ and that claim (3) is true for every $\eta<\alpha$. If $\alpha$ is a successor ordinal, then $x_{\alpha}=f(x_{\alpha-1})$, and thus, for $0\leq\beta<\eta\leq\alpha$, we have
    $x_{\beta}\,\mathcal{C}\,x_{\alpha-1}\implies x_{\beta}\,\mathcal{C}\,x_{\alpha}.$
    Assume that $\alpha$ is a limit ordinal and take $\beta<\alpha$ and $\epsilon>0$. Since $x_{\alpha}\in S^{*}_{\alpha}(x)$, there is $\beta<\eta<\alpha$ such that $d(x_{\eta},x_{\alpha})<\frac{\epsilon}{2}$. By the induction hypothesis, we have $x_{\beta}\,\mathcal{C}_{\frac{\epsilon}{2}}\,x_{\eta}$, so by the triangle inequality $x_{\beta}\,\mathcal{C}_{\epsilon}\,x_{\alpha}$, and since $\epsilon$ was arbitrary, we have $x_{\beta}\,\mathcal{C}\,x_{\alpha}$.
  • Step 3
    Now, if for every $\eta<\alpha$ we have $x_{\eta}\notin S_{\eta}(x)$, then $x_{\beta}\neq x_{\gamma}$ for every distinct $\beta,\gamma<\alpha$, so that the application
    $\eta\mapsto x_{\eta}\quad\quad(0\leq\eta<\alpha)$ (4)
    is a bijection between $\alpha=\{\eta:0\leq\eta<\alpha\}$ and $S_{\alpha}(x)$, so we have $|S_{\alpha}(x)|=|\alpha|$. Assuming $x_{\alpha}\notin S_{\alpha}(x)$ for every ordinal $\alpha$, by Hartogs’ Lemma ([14]) there would be an ordinal $\lambda$ so large that $|S_{\lambda}(x)|>|X|$, in the sense that there is no injection from $S_{\lambda}(x)$ into $X$. But this is impossible since $S_{\alpha}(x)\subseteq X$. Therefore, $x_{\alpha}\in S_{\alpha}(x)$ for some ordinal $\alpha$, so that:
    $x_{\alpha}=x_{\beta}\text{ for some }\beta<\alpha.$ (5)
    This implies, using claim (3) with $\eta=\alpha$, that $x_{\alpha}$ is chain-recurrent.
Remark 3.2.
Notice that, since in Step 2 we built every chain using only points indexed by ordinals, for every $\beta<\eta\leq\alpha$ and every $\epsilon>0$, there is an $\epsilon$-chain from $x_{\beta}$ to $x_{\eta}$ whose points belong to $S_{\beta,\eta+1}(x)$.
Remark 3.3.
The argument in the previous proof does not use the Axiom of Choice. In particular, Hartogs’ Lemma holds in ZF ([15], Chapter 1). The same holds for the other existence results in this work.
Since a compact space has at most cardinality continuum (see [16, Corollary 3.1.30]), assuming the Generalized Continuum Hypothesis we can take, in the previous proof, $\lambda\leq\aleph_{2}=2^{\mathfrak{c}}$.
If we replace the chain relation $\mathcal{C}$ with the relation $\widetilde{\mathcal{N}}$, the argument in the previous proof is no longer applicable. In fact, in Step 2, for $\alpha=1$ claim (3) is also true by replacing $\mathcal{C}$ with $\widetilde{\mathcal{N}}$, so we can assume as inductive hypothesis that, for $\alpha>1$, claim (3) is true, with $\widetilde{\mathcal{N}}$ instead of $\mathcal{C}$, for every $\eta<\alpha$. However, when $\alpha$ is a successor ordinal, even if we have $x_{\beta}\,\widetilde{\mathcal{N}}\,x_{\alpha-1}$ for $0\leq\beta<\eta\leq\alpha$, it is not true, in general, that $x_{\beta}\,\widetilde{\mathcal{N}}\,f(x_{\alpha-1})$ when $f$ is discontinuous at $x_{\alpha-1}$. This of course is consistent with the example given in Eq. (1) in the Introduction.
The following result is an easy consequence of the previous theorem.
Theorem 3.4.
Let $(X,f)$ be a compact dynamical system. Then for every $x\in X$ there is $y\in X$ such that $y$ is chain-recurrent and $x\,\mathcal{C}\,y$.
Proof.
In the previous argument, and specifically in Eq.(5), take $y=x_{\alpha}$. Since $x=x_{0}$ was arbitrary, the result follows. ∎
Notice that, when the map is continuous, the result above can be strengthened by replacing the last relation $x\,\mathcal{C}\,y$ by $x\,\mathcal{R}\,y$, and it is a consequence of the existence of a complete Lyapunov function (a seminal work is [1]).
Theorem 3.5.
Let $(X,f)$ be a compact dynamical system. Then $GR(f)\neq\emptyset$.
Proof.
Again we can assume that every orbit in $X$ is infinite (otherwise the result follows immediately). Let $\mathcal{D}$ be the set of all metrics compatible with the topology on $X$, and pick $d\in\mathcal{D}$. Pick $x\in X$ and set $x_{0}:=x$. For $k\in\mathbb{N}$, set $x_{k}:=f(x_{k-1})$. For any ordinal $\alpha$, we define the sets $S_{\alpha}(x)$ and $S^{*}_{\alpha}(x)$ as done in Step 1 of the proof of Theorem 3.1. Notice that this construction depends only on the topology and on the map $f$, but does not depend on the metric $d$.
Following the steps in the proof of Theorem 3.1, we want to prove by induction on $\alpha$ that
$x_{\beta}\,\mathcal{SC}_{d}\,x_{\eta}\quad\text{ whenever }\quad 0\leq\beta<\eta\leq\alpha.$ (6)
If $\alpha=1$, the claim is obviously true; therefore, we assume that claim (6) is true for every $\eta<\alpha$. If $\alpha$ is a successor ordinal, then $x_{\alpha}=f(x_{\alpha-1})$, and thus, for $0\leq\beta<\eta\leq\alpha$, we have
$x_{\beta}\,\mathcal{SC}_{d}\,x_{\alpha-1}\implies x_{\beta}\,\mathcal{SC}_{d}\,x_{\alpha}.$
Assume that $\alpha$ is a limit ordinal and take $\beta<\alpha$ and $\epsilon>0$. Since $x_{\alpha}\in S^{*}_{\alpha}(x)$, there is $\beta<\eta<\alpha$ such that $d(x_{\eta},x_{\alpha})<\frac{\epsilon}{2}$. By the induction hypothesis, there is a strong $(\frac{\epsilon}{2},d)$-chain $x_{\beta}=x_{0},x_{1},\ldots,x_{n}=x_{\eta}$ from $x_{\beta}$ to $x_{\eta}$. By the triangle inequality, we have:
$\sum_{i=0}^{n-2}d(f(x_{i}),x_{i+1})+d(f(x_{n-1}),x_{\alpha})\leq\sum_{i=0}^{n-1}d(f(x_{i}),x_{i+1})+d(x_{\eta},x_{\alpha})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.$
Therefore, the points $x_{\beta}=x_{0},x_{1},\ldots,x_{n-1},x_{\alpha}$ form a strong $(\epsilon,d)$-chain from $x_{\beta}$ to $x_{\alpha}$. Since $\epsilon$ was arbitrary, it follows that $x_{\beta}\,\mathcal{SC}_{d}\,x_{\alpha}$. Then, by an identical argument as in Step 3 above (with (6) in place of (3)), we have that there is some ordinal $\alpha$ such that $x_{\alpha}\in\mathcal{SCR}_{d}$.
Since claim (6) is valid for every choice of the metric $d$ (compatible with the topology), we have
$x_{\alpha}\in\bigcap_{d\in\mathcal{D}}\mathcal{SCR}_{d}(f)=GR(f).$
We prove now that every compact system has a chain-transitive subsystem. For this we need a preliminary lemma, which deepens our understanding of the construction in the proof of Theorem 3.1.
Lemma 3.6.
Let $(X,f)$ be a compact dynamical system. Then $X$ has an invariant, $\mathcal{C}$-transitive subset.
Proof.
If some point is periodic, the periodic orbit is clearly an invariant, $\mathcal{C}$-transitive subset of $X$. We can then assume that every orbit is infinite. We will use the notation defined in the proof of Theorem 3.1. For every $\alpha$ limit ordinal and $\beta\leq\alpha$ we set
$\Gamma_{\beta,\alpha}:=\{\beta\}\cup\{\beta<\eta<\alpha:\text{$\eta$ is a limit ordinal}\}.$
Pick $x\in X$. Let $\lambda$ be an ordinal such that $|\lambda|>|X|$ and $S_{\lambda}(x)$ a set defined by means of the transfinite procedure described in Step 1 of the proof of Theorem 3.1. Let $\alpha\leq\lambda$ be the least ordinal greater than 0 such that $x_{\alpha}\in S_{\alpha}(x)$. Then there is an ordinal $0\leq\beta<\alpha$ such that $x_{\alpha}=x_{\beta}$ and $x_{\alpha}\in CR_{f}$.
If $\alpha$ is a limit ordinal, then the set $S_{\beta,\alpha}(x)$ can be written as:
$S_{\beta,\alpha}(x)=\bigcup_{\eta\in\Gamma_{\beta,\alpha}}\mathcal{O}^{0}(x_{\eta}),$
that is, it is a union of orbits and therefore is invariant. Analogously, if $\alpha$ is a successor ordinal, that is $\alpha=\gamma+k$ for some limit ordinal $\gamma\geq\beta$ and $k\in\mathbb{N}$ the set $S_{\beta,\alpha}(x)$ can be written as:
$S_{\beta,\alpha}(x)=\bigcup_{\eta\in\Gamma_{\beta,\gamma}}\mathcal{O}^{0}(x_{\eta})\cup\bigcup_{h=0}^{k-1}f^{h}(x_{\gamma}),$
which is again invariant, since in this case $f^{k}(x_{\gamma})=x_{\alpha}=x_{\beta}$. In both cases, since claim (3) holds, that is, $x_{\beta}\,\mathcal{C}\,x_{\eta}$ for every $0\leq\beta<\eta\leq\alpha$ and $x_{\alpha}=x_{\beta}$, by the transitivity of $\mathcal{C}$, it follows that $S_{\beta,\alpha}(x)$ is $\mathcal{C}$-transitive. ∎
Remark 3.7.
Notice that by Remark 3.2 and since $x_{\alpha}=x_{\beta}$ it follows that for every $z,y\in S_{\beta,\alpha}(x)$ and for every $\epsilon>0$ there exists an $\epsilon$-chain from $z$ to $y$ whose points belong to $S_{\beta,\alpha}(x)$. This is equivalent to say that the set $S_{\beta,\alpha}(x)$ is internally chain transitive.
Remark 3.8.
Notice that by claim (6) and since $x_{\alpha}=x_{\beta}$, by the transitivity of $\mathcal{SC}_{d}$, it follows that $S_{\beta,\alpha}(x)$ is $\mathcal{SC}_{d}$-transitive.
Theorem 3.9.
Let $(X,f)$ be a compact dynamical system. Then $(X,f)$ has a $\mathcal{C}$-transitive subsystem.
Proof.
Let us first of all observe that, to prove the statement, we need to show that there is a closed, invariant and internally chain transitive subset of $X$. In particular, the last property ensures that we obtained an object which is a chain-transitive subsystem of $(X,f)$, and not just a closed, invariant and chain-transitive subset. We will use the notation defined in the proof of Theorem 3.1: for $x\in X$ and for ordinals $\zeta<\iota$, let $S_{\zeta,\iota}(x)$ and $S_{\zeta}(x)=S_{0,\zeta}(x)$ be sets defined as in Step 1 of the proof of Theorem 3.1.
Suppose, towards a contradiction, that $(X,f)$ does not have a chain-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By Theorem 3.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 3.6 and Remark 3.7, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally chain transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. By Remark 3.2 and by the triangle inequality, we have that for every $x\in S_{\alpha_{0}}(x^{0})$ and for every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. If $y\neq x^{0}$, then we set $x^{1}:=y$. Otherwise, if $y=x^{0}$, it follows that $S_{\alpha_{0}}(x^{0})$ is an invariant, internally chain transitive subset. Since it cannot be a closed set, there is some $z\in(S_{\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\alpha_{0}}(x^{0})$, and we set $x^{1}:=z\,(\neq x^{0})$. We then have $S_{\alpha_{0}}(x^{0})\,\mathcal{C}\,x^{1}$, and for every $x\in S_{\alpha_{0}}(x^{0})$ and for every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $x^{1}$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{x^{1}\}$.
Summarizing, we have shown that the set $\{x^{0},x^{1}\}$ has the following properties:
  • (a1)
    there exists a least ordinal $\alpha_{0}$ such that $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$ and $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally chain transitive subset;
  • (b1)
    $S_{\alpha_{0}}(x^{0})\,\mathcal{C}\,x^{1}$, and for every $x\in S_{\alpha_{0}}(x^{0})$ and for every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $x^{1}$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{x^{1}\}$;
  • (c1)
    $x^{0}\neq x^{1}$.
Let us proceed now by transfinite induction. Let $\lambda>1$ be an ordinal number. Assume that $x^{\gamma}$ has been defined for every $0\leq\gamma<\lambda$ and that the set $\{x^{\gamma}\}_{\gamma<\lambda}$ has the following property:
  1. (aλ)
    for every $0\leq\gamma<\lambda$, there exists a least ordinal $\alpha_{\gamma}$ such that $x^{\gamma}_{\alpha_{\gamma}}=x^{\gamma}_{\beta_{\gamma}}$ for some $\beta_{\gamma}<\alpha_{\gamma}$ and $S_{\beta_{\gamma},\alpha_{\gamma}}(x^{\gamma})$ is an invariant, internally chain transitive subset.
Moreover, setting for every $\beta<\alpha\leq\lambda$
$M_{\beta,\alpha}:=\bigcup_{\beta\leq\gamma<\alpha}S_{\alpha_{\gamma}}(x^{\gamma}),$
assume also the following properties:
  1. (bλ)
    For every $\eta<\xi<\lambda$, we have $S_{\alpha_{\eta}}(x^{\eta})\,\mathcal{C}\,x^{\xi}$, and for every $x\in S_{\alpha_{\eta}}(x^{\eta})$ and every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $x^{\xi}$ whose points belong to $M_{\eta,\xi}\cup\{x^{\xi}\}$;
  2. (cλ)
    $x^{\eta}\neq x^{\xi}$ whenever $\eta\neq\xi$ with $\eta,\xi<\lambda$
We say that a point $z$ verifies property Pλ if, for every $\eta<\lambda$, we have $S_{\alpha_{\eta}}(x^{\eta})\,\mathcal{C}\,z$, and for every $x\in S_{\alpha_{\eta}}(x^{\eta})$ and every $\epsilon>0$, there exists an $\epsilon$-chain from $x$ to $z$ whose points belong to $M_{\eta,\lambda}\cup\{z\}$.
Notice that, assuming (bλ), to prove property (bλ+1) it is sufficient to show that the point $x^{\lambda}$ verifies property Pλ. We now proceed with the definition of the point $x^{\lambda}$.
  • Case 1.
    $\lambda$ is a successor ordinal.
    Consider the set $S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1})$, which is invariant and internally chain transitive by property (aλ). Since it cannot be a closed set, there exists some
    $y\in(S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1}))^{\prime}\setminus S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1}).$
    The point $y$ verifies property Pλ. Indeed, by Remark 3.2, and by the triangle inequality, we have that for every $x\in S_{\alpha_{\lambda-1}}(x^{\lambda-1})$ and for every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{\lambda-1}}(x^{\lambda-1})\cup\{y\}$. Fix $\eta<\lambda$ and $\epsilon>0$ and take $x\in S_{\alpha_{\eta}}(x^{\eta})$. Then there is an $\epsilon$-chain from $x^{\lambda-1}\in S_{\alpha_{\lambda-1}}(x^{\lambda-1})$ to $y$ whose points belong to $S_{\alpha_{\lambda-1}}(x^{\lambda-1})\cup\{y\}$. By property (bλ), for every $\epsilon>0$ there exists an $\epsilon$-chain from $x$ to $x^{\lambda-1}$ whose points belong to $M_{\eta,\lambda-1}\cup\{x^{\lambda-1}\}$. Therefore, noting that $M_{\eta,\lambda}=M_{\eta,\lambda-1}\cup S_{\alpha_{\lambda-1}}(x^{\lambda-1})$ and using the transitivity of $\mathcal{C}$, we have that there is an $\epsilon$-chain from $x$ to $y$ whose points belong to $M_{\eta,\lambda}\cup\{y\}$.
    Consider the following two cases.
    • If $y\notin\{x^{\gamma}\}_{\gamma<\lambda}$, then we set $x^{\lambda}:=y$. Then $x^{\lambda}$ verifies property Pλ and (bλ+1) is satisfied.
    • Assume that $y=x^{\xi_{0}}$ for some $\xi_{0}<\lambda$. By property (bλ) and since $y$ verifies property Pλ, we have that the set $M_{\xi_{0},\lambda}$ is invariant and internally chain transitive. Since this set cannot be closed, there is some $z_{1}\in(M_{\xi_{0},\lambda})^{\prime}\setminus M_{\xi_{0},\lambda}$, so that $z_{1}\notin\{x^{\gamma}\}_{\xi_{0}\leq\gamma<\lambda}$. Since $M_{\xi_{0},\lambda}$ is internally chain transitive and $z_{1}\in(M_{\xi_{0},\lambda})^{\prime}$, by the triangle inequality it follows that for every $x\in M_{\xi_{0},\lambda}$ there exists an $\epsilon$-chain from $x$ to $z_{1}$ whose points belong to $M_{\xi_{0},\lambda}\cup\{z_{1}\}$. By property (bλ) and using the transitivity of $\mathcal{C}$, we have that the point $z_{1}$ verifies property Pλ.
      If $z_{1}=x^{\xi_{1}}$ for some $\xi_{1}<\xi_{0}$, we repeat the same argument replacing $\xi_{1}$ with $\xi_{0}$ and observing that the set $M_{\xi_{1},\lambda}$ is invariant and internally chain transitive.
      Since there is no infinitely decreasing sequence of ordinals, we can repeat the previous construction up to a certain $k\in\mathbb{N}$, after which we must have $z_{k+1}\notin\{x^{\gamma}\}_{\gamma<\lambda}$. Then we set $x^{\lambda}:=z_{k+1}$ and since it verifies property Pλ, property (bλ+1) is satisfied.
[…]
  • Case 2.
    $\lambda$ is a limit ordinal.
    Since |$|\{x^{\xi}\ |\ \gamma\leq\xi<\lambda\}|\geq\aleph_{0}$ for every $\gamma<\lambda$, by compactness (see [13, Th. 26.9, p. 169]) we can pick
    $y\in\bigcap_{\gamma<\lambda}(\{x^{\xi}\ |\ \gamma\leq\xi<\lambda\})^{\prime}.$
    The point $y$ verifies property Pλ. Indeed, fix $\eta<\lambda$ and $\epsilon>0$ and take $x\in S_{\alpha_{\eta}}(x^{\eta})$. Then there exists $\eta<\xi<\lambda$ such that $d(x^{\xi},y)<\frac{\epsilon}{2}$. By property (bλ), there exists an $\frac{\epsilon}{2}$-chain from $x$ to $x^{\xi}$ whose points belong to $M_{\eta,\xi}\cup\{x^{\xi}\}$. By the triangle inequality, it follows that there exists an $\epsilon$-chain from $x$ to $y$ whose points belong to $M_{\eta,\xi}\cup\{y\}\subseteq M_{\eta,\lambda}\cup\{y\}$.
    Consider the following two cases.
    • If $y\notin\{x^{\gamma}\}_{\gamma<\lambda}$, then we set $x^{\lambda}:=y$. Then, $x^{\lambda}$ verifies property Pλ and (bλ+1) is satisfied.
    • Assume that $y=x^{\xi_{0}}$ for some $\xi_{0}<\lambda$. By property (bλ) and since $y$ verifies property Pλ we have that the set $M_{\xi_{0},\lambda}$ is invariant and internally chain transitive. Since this set cannot be closed, there exists $z_{1}$ such that $z_{1}\in\left(M_{\xi_{0},\lambda}\right)^{\prime}\setminus M_{\xi_{0},\lambda}.$ Therefore, we have that $z_{1}\notin\{x^{\gamma}\}_{\xi_{0}\leq\gamma<\lambda}$. Since $M_{\xi_{0},\lambda}$ is internally chain transitive and $z\in(M_{\xi_{0},\lambda})^{\prime}$, by property (bλ) and by the triangle inequality it follows that for every $x\in M_{\xi_{0},\lambda}$ there exists an $\epsilon$-chain from $x$ to $z_{1}$ whose points belong to $M_{\xi_{0},\lambda}\cup\{z_{1}\}$. By property (bλ) and using the transitivity of $\mathcal{C}$, we have that the point $z_{1}$ verifies property Pλ.
      If $z_{1}=x^{\xi_{1}}$ for some $\xi_{1}<\xi_{0}$, we repeat the same argument replacing $\xi_{1}$ with $\xi_{0}$ and observing that the set $M_{\xi_{1},\lambda}$ is invariant and internally chain transitive.
      Since there is no infinitely decreasing sequence of ordinals, we can repeat the previous construction up to a certain $k\in\mathbb{N}$, after which we must have $z_{k+1}\notin\{x^{\gamma}\}_{\gamma<\lambda}$. Then we set $x^{\lambda}:=z_{k+1}$ and since it verifies property Pλ, property (bλ+1) is satisfied.
[…]
Assuming that $(X,f)$ does not have a chain-transitive subsystem, we find that the application
$\gamma\mapsto x^{\gamma}\quad(0\leq\gamma<\lambda)$
is a bijection between $\lambda$ and the set $\{x^{\gamma}\}_{0\leq\gamma<\lambda}$, so we have that |$|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|=|\lambda|$. By Hartogs’ Lemma ([14]), we can take $\lambda$ so large that |$|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|>|X|$, which is a contradiction.
Therefore, there exists an ordinal $\nu<\lambda$ such that it is impossible to define the point $x^{\nu}$. More precisely, we have that, if $\nu$ is a successor ordinal, one cannot define the point $x^{\nu}$ if one of the following verifies:
  • The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is closed, so that $(S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1}))^{\prime}\subseteq S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$.
  • The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is not closed and there exists an ordinal $\xi<\nu$ such that the set $M_{\xi,\nu}$ is a closed, invariant and internally chain transitive subset, that is a chain-transitive subsystem.
On the other hand, if $\nu$ is a limit ordinal, then we cannot define the point $x^{\nu}$ if for every $y\in\bigcap_{\gamma<\nu}(\{x^{\xi}\ |\ \gamma\leq\xi<\nu\})^{\prime}$, we have that $y\in\{x^{\gamma}\}_{\gamma<\nu}$. This implies that there exists an ordinal $\xi<\nu$ such that $M_{\xi,\nu}$ is a closed, invariant and internally chain transitive subset, that is a chain-transitive subsystem. ∎
Figure 2: Graphic representation of (a sample case of) the reasoning in the proof of Theorem 3.9.
Notice that, assuming $f$ continuous, we can strengthen the previous result replacing “chain-transitive” by “minimal” (see [2], Proposition 2.20).
From the previous theorem we can readily deduce the following result.
Theorem 3.10.
Let $X$ be a compact dynamical system. Then every $x\in X$ is in chain-recurrence relation with every point of certain closed, invariant, internally chain transitive subset $S\subseteq X$.
Proof.
Pick any $x^{0}\in X$. By Theorem 3.9 there exists a least ordinal $\nu$ such that the set $\{x^{\gamma}\}_{\gamma<\nu}$ is defined, while it is impossible to define the point $x^{\nu}$.
If $\nu$ is a successor ordinal, then one of the following verifies:
  1. 1.
    The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is a closed, invariant and internally chain transitive subset. By construction of the set $\{x^{\gamma}\}_{\gamma<\nu}$ we have $x^{0}\,\mathcal{C}\,S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ and the claim is verified with $x=x^{0}$ and $S=S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$.
  2. 2.
    The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is not closed and there exists an ordinal $\xi<\nu$ such that the set $M_{\xi,\nu}$ is closed, invariant and internally chain transitive. Again, by construction, we have $x^{0}\,\mathcal{C}\,M_{\xi,\nu}$ and the claim is verified with $x=x^{0}$ and $S=M_{\xi,\nu}$.
On the other hand, if $\nu$ is a limit ordinal and we cannot define the point $x^{\nu}$, then there exists an ordinal $\xi<\nu$ such that $M_{\xi,\nu}$ is a closed, invariant and internally chain transitive subset. Moreover, we have $x^{0}\,\mathcal{C}\,M_{\xi,\nu}$ and the claim is verified with $x=x^{0}$ and $S=M_{\xi,\nu}$.
Since $x^{0}$ was arbitrary, the result follows. ∎
Similarly to what we observed after Theorem 3.4, also here, if the map is continuous, one has a stronger result, as we can replace, in the previous statement, “chain-recurrence relation” by “recurrence relation” (see for instance, Lemma 2.1 in [9]).
4 Nested chains
Looking closer at the proof of Theorem 3.1, we can observe that in fact something stronger can be said. Indeed, as we will prove (see Theorem 4.6), we obtained a point $x$ belonging to the diagonal of a relation $\mathcal{C}_{{}_{\subseteq}}$ on $X^{2}$, verifying in general $\mathcal{C}_{{}_{\subseteq}}\subseteq\mathcal{C}$, defined as follows:
Definition 4.1.
Let $(X,f)$ be a compact dynamical system and $d:X^{2}\to\mathbb{R}_{0}^{+}$ be a metric compatible with the topology on $X$. We say that $x\,\mathcal{C}_{{}_{\subseteq}}y$ if, for a non-increasing sequence of positive reals $\{\epsilon_{n}\}_{n\in\mathbb{N}}$ converging to zero, there is a collection $\{C_{\epsilon_{n}}\}_{n\in\mathbb{N}}$ of $\epsilon_{n}$-chains from $x$ to $y$ such that $C_{\epsilon_{n}}\subseteq C_{\epsilon_{n+1}}$ for every $n\in\mathbb{N}$. We call $\mathcal{C}_{{}_{\subseteq}}$ the nested-chain relation and $\{C_{\epsilon_{n}}\}_{n\in\mathbb{N}}$ the nested $\epsilon$-chains.
A point $x\in X$ is called $\mathcal{C}_{\subseteq}$-recurrent if $x\,\mathcal{C}_{\subseteq}\,x$.
It is easy to see that, if $x\,\mathcal{C}_{{}_{\subseteq}}y$, then the sequence of nested $\epsilon$-chains $\{C_{\epsilon_{n}}\}_{n\in\mathbb{N}}$ works (possibly up to subsequence) for every positive sequence $\{\epsilon_{n}\}_{n\in\mathbb{N}}$ converging to zero, in the sense that, if $\delta_{j}\to 0$, then there is a subsequence $\{n_{j}\}$ such that, for every $j$, $C_{\epsilon_{n_{j}}}$ is a $\delta_{j}$ chain from $x$ to $y$ and we have $C_{\epsilon_{n_{j}}}\subseteq C_{\epsilon_{n_{j+1}}}$. The relation $\mathcal{C}_{\subseteq}$ can be formulated, just like $\mathcal{C}$, in pure topological terms:
Theorem 4.2.
Let $(X,f)$ be a compact dynamical system. Then, $x\,\mathcal{C}_{\subseteq}\,y$ if and only if, for a sequence of open neighborhood $\{U_{i}\}_{i\in\mathbb{N}}$ of $\Delta X$ such that:
  • 1)
    $\overline{U_{i+1}}\subseteq U_{i}$ for every $i\in\mathbb{N}$,
  • 2)
    $\bigcap_{i\in\mathbb{N}}\overline{U_{i}}=\Delta X$,
for every $i\in\mathbb{N}$ there exist $n_{i}\in\mathbb{N}$ and a set $C_{i}=\{x=x_{0},\dots,x_{n_{i}}=y\}$ such that:
  • 3)
    $(f(x_{j}),x_{j+1})\in U_{i}\text{ for every }j=0,\ldots,n_{i}-1$,
and the family $\{C_{i}\}_{i\in\mathbb{N}}$ verifies:
  • 4)
    $C_{i}\subseteq C_{i+1}$ for every $i\in\mathbb{N}$.
Proof.
Let $d$ be a metric compatible with the topology on $X$. Let us assume that $x\,\mathcal{C}_{\subseteq}\,y$. Therefore, there are a non-increasing sequence of positive reals $\{\epsilon_{k}\}_{k\in\mathbb{N}}$ converging to zero and a collection $\{C_{\epsilon_{k}}\}_{k\in\mathbb{N}}$ of $\epsilon_{k}$-chains from $x$ to $y$ such that $C_{\epsilon_{k}}\subseteq C_{\epsilon_{k+1}}$ for every $k\in\mathbb{N}$. Then, there exists a strictly decreasing subsequence $\{\epsilon_{k_{i}}\}_{i\in\mathbb{N}}$ of $\{\epsilon_{k}\}_{k\in\mathbb{N}}$. For every $i\in\mathbb{N}$, we set $U_{i}:=B_{\epsilon_{k_{i}}}(\Delta X)$. Since $\epsilon_{k_{i}}\to 0$ as $i\to\infty$, we have $\bigcap_{i\in\mathbb{N}}\overline{U_{k_{i}}}=\Delta X$ and since the sequence $\{\epsilon_{k_{i}}\}_{i\in\mathbb{N}}$ is strictly decreasing, it follows that $\overline{U_{i+1}}\subseteq U_{i}$ for every $i\in\mathbb{N}$. Set $C_{i}:=C_{\epsilon_{k_{i}}}$ for every $i\in\mathbb{N}$. Then
$C_{i}=C_{\epsilon_{k_{i}}}\subseteq C_{\epsilon_{k_{i+1}}}=C_{i+1}\quad\text{for every }i\in\mathbb{N}.$
Moreover, since $C_{\epsilon_{k_{i}}}$ is an $\epsilon_{k_{i}}$-chain from $x$ to $y$ we have $C_{i}=C_{\epsilon_{k_{i}}}=\{x=x_{0},\ldots,x_{n_{i}}=y\}$ for some $n_{i}\in\mathbb{N}$ and
$(f(x_{j}),x_{j+1})\in B_{\epsilon_{k_{i}}}(\Delta X)=U_{i}\quad\text{for every}j=0,\ldots,n_{i}-1.$
Let us prove the converse implication. Suppose that for $x,y\in X$ there are a sequence of open neighborhoods $\{U_{i}\}_{i\in\mathbb{N}}$ of $\Delta X$ with properties 1) and 2) and a collection of sets $\{C_{i}\}_{i\in\mathbb{N}}$ with properties 3) and 4). Since $X$ is compact and exploiting the properties of $\{U_{i}\}_{i\in\mathbb{N}}$ (see [2, Prop. A.28]), we can define the sequence $\{\epsilon_{i}\}_{i\in\mathbb{N}}$ by induction as follows.
Set $k_{0}:=1$. By [2, Prop. A.28-(2)] there is $\epsilon_{1}>0$ such that $B_{\epsilon_{1}}(\Delta X)\subseteq U_{k_{0}}$. By [2, Prop. A.28-(3)] there exists $k_{1}>k_{0}$ such that $U_{k_{1}}\subseteq\overline{U_{k_{1}}}\subseteq B_{\epsilon_{1}}(\Delta X)$. Assume that $\epsilon_{i}$ and $k_{i}$ have been defined for every $0\leq i\leq n$ such that $U_{k_{i}}\subseteq B_{\epsilon_{i}}(\Delta X)\subseteq U_{k_{i-1}}$. By [2, Prop. A.28-(2)] there exists $\epsilon_{n+1}>0$ such that $B_{\epsilon_{n+1}}(\Delta X)\subseteq U_{k_{n}}$. By [2, Prop. A.28-(3)] there exists $k_{n+1}$ sufficiently large that $U_{k_{n+1}}\subseteq\overline{U_{k_{n+1}}}\subseteq B_{\epsilon_{n+1}}(\Delta X)$.
By construction, the sequence $\{\epsilon_{i}\}_{i\in\mathbb{N}}$ is non-increasing and by 2) it follows that $\epsilon_{i}\to 0$ as $i\to\infty$. Set $C_{\epsilon_{i}}:=C_{k_{i}}$ for every $i\in\mathbb{N}$. Since $U_{k_{i}}\subseteq B_{\epsilon_{i}}(\Delta X)$, by properties 3) and 4), it follows that $C_{\epsilon_{i}}$ is an $\epsilon_{i}$-chain from $x$ to $y$ for every $i\in\mathbb{N}$, and $\{C_{\epsilon_{i}}\}_{i\in\mathbb{N}}$ is a collection of nested $\epsilon$-chains. We conclude that $x\,\mathcal{C}_{\subseteq}\,y$. ∎
In general, $\mathcal{C}_{\subseteq}$ is stronger than $\mathcal{C}$. Let us show a continuous example in which $\mathcal{C}_{\subseteq}\subsetneq\mathcal{C}$.
Example 4.3.
Let $\{0,1\}^{\omega}$ be the set of infinite binary words and $\sigma:\{0,1\}^{\omega}\rightarrow\{0,1\}^{\omega}$ the full shift defined by
$\left(\sigma(u)\right)_{n}=u_{n+1},\quad\text{for all }n\in\mathbb{N},$
where by $u_{k}$ we indicate the digit indexed by the positive integer $k$ of the word $u$. Set
$d(u,v)=2^{-(n-1)},\ \text{where $n:=\min\{n\in\mathbb{N}:u_{n}\neq v_{n}\}$}.$
Let $\Sigma\subseteq\{0,1\}^{\omega}$ be the subshift defined by the following countable set of forbidden words:
$\{01^{j}0\text{ such that }j\in\mathbb{N}\}\quad\cup\quad\{1^{j}0^{h}\text{ such that }j>h\}.$
Defining the words:
$w_{k}:=1^{k}0^{k}1^{\infty},$
the subshift $\Sigma$ consists of the words
$\sigma^{j}(w_{k})\quad(j\in\mathbb{N}_{0},\ k\in\mathbb{N})$
and the limit points
$\lim_{k\to\infty}w_{k}=1^{\infty}\quad,\quad\lim_{k\to\infty}\sigma^{k}w_{k}=0^{\infty}.$
Consider the subshift $\Sigma$. We want to show that $(1^{\infty},0^{\infty})\in\mathcal{C}\setminus\mathcal{C}_{\subseteq}$.
Take $\epsilon>0$, and let $n>0$ be such that $2^{-n}<\epsilon$. If we set $x_{i}:=\sigma^{(i-1)}(w_{n})$ for $i=1,\ldots,n$, then the points $1^{\infty}=x_{0},x_{1},\ldots,x_{n},x_{n+1}=0^{\infty}$ form an $\epsilon$-chain from $1^{\infty}$ to $0^{\infty}$. Indeed, we have that
  • $d(\sigma(1^{\infty}),x_{1})=d(1^{\infty},1^{n}0^{n}1^{\infty})=2^{-n}<\epsilon$;
  • $d(\sigma(x_{i}),x_{i+1})=d(x_{i+1},x_{i+1})=0$ for $i=1,\ldots,n-1$;
  • $d(\sigma(x_{n}),0^{\infty})=d(\sigma^{n}(w_{n}),0^{\infty})=d(0^{n}1^{\infty},0^{\infty})=2^{-n}<\epsilon$.
[…]
$\mathcal{O}(x_{1})=\{1^{k-j}0^{p}1^{\infty}\}_{j=1}^{k-1}\cup\{0^{p-j}1^{\infty}\}_{j=0}^{p-1}\cup\{1^{\infty}\}.$
Since $p\geq k\geq 2m+1$, we have that $11^{h}0^{m}$ 1 is not a subword of any element of $\mathcal{O}(x_{1})$, thus we have:
$d(u,\mathcal{O}(x_{1}))\geq 2^{-2m}>\epsilon_{2m}.$
By the definition of $\Sigma$, it is clear that $(x,u)\notin\mathcal{C}_{\epsilon_{2m}}$ for all $x\in\mathcal{O}(x_{1})$. Then $u\notin C_{\epsilon_{2m}}$ and we can conclude.
Remark 4.4.
[…]
Let us now prove that the chain-recurrent point whose existence is proved in Theorem 3.1 is, in fact, a $\mathcal{C}_{\subseteq}$-recurrent point. For this, we start with the following lemma showing that claim (3) also applies with respect to the relation $\mathcal{C}_{\subseteq}$.
Lemma 4.5.
Let $(X,f)$ be a compact dynamical system and $x\in X$. For ordinals $\zeta<\lambda$, let $S_{\zeta,\lambda}(x)$ and $S_{\lambda}(x)=S_{0,\lambda}$ be sets defined as in Step 1 of the proof of Theorem 3.1. Then
$x_{\eta}\,\mathcal{C}_{\subseteq}\,x_{\xi}\quad\text{whenever}\quad 0\leq\eta<\xi\leq\lambda.$
Proof.
Let $d$ be a metric compatible with the topology on $X$ and take $0\leq\eta<\xi\leq\lambda$. Let $\{\epsilon_{m}\}_{m\in\mathbb{N}}$ be a non-increasing sequence of positive reals converging to zero. By Theorem 3.1 and Remark 3.2 there is an $\epsilon_{1}$-chain $C_{\epsilon_{1}}\subseteq S_{\eta,\xi+1}(x)$ from $x_{\eta}$ to $x_{\xi}$.
Proceeding by induction, assume that, for every positive integer $m\leq n$, there is an $\epsilon_{m}$-chain $C_{\epsilon_{m}}\subseteq S_{\eta,\xi+1}(x)$ from $x_{\eta}$ to $x_{\xi}$, and that, for all $m<n$, one has $C_{\epsilon_{m}}\subseteq C_{\epsilon_{m+1}}$. Assume that $C_{\epsilon_{n}}$ consists of the points:
$x_{\eta}=x_{\beta_{0}},x_{\beta_{1}},\dots,x_{\beta_{k}}=x_{\xi}\quad\quad(\eta=\beta_{0}<\beta_{1}<\dots<\beta_{k}=\xi\leq\lambda).$
By Remark 3.2, for every $0\leq j<k$, there is an $\epsilon_{n+1}$-chain $C_{\epsilon_{n+1}}^{j}\subseteq S_{\beta_{j},(\beta_{j+1}+1)}(x)$ from $x_{\beta_{j}}$ to $x_{\beta_{j+1}}$, consisting of the points:
$x_{\beta_{j}}=x_{\beta_{j_{1}}},x_{\beta_{j_{2}}},\dots,x_{\beta_{j_{h_{j}}}}=x_{\beta_{j+1}}\quad\quad(\beta_{j}=\beta_{j_{1}}<\beta_{j_{2}}<\dots<\beta_{j_{h_{j}}}=\beta_{j+1}).$
Therefore, the points
$x_{\eta}=x_{\beta_{0}}=x_{{\beta_{0}}_{1}},x_{{\beta_{0}}_{2}},\dots,x_{{\beta_{0}}_{h_{0}}}=x_{{\beta_{1}}_{1}},\dots,x_{{\beta_{1}}_{h_{1}}}=x_{{\beta_{2}}_{1}},\dots\dots\dots,x_{{\beta_{k-1}}_{h_{k-1}}}=x_{\beta_{k}}=x_{\xi}$
form an $\epsilon_{n+1}$-chain $C_{\epsilon_{n+1}}$ from $x_{\eta}$ to $x_{\xi}$ such that
$C_{\epsilon_{n}}\subseteq C_{\epsilon_{n+1}}=\bigcup_{j=0}^{k-1}C_{\epsilon_{n+1}}^{j}\subseteq S_{\eta,\xi+1}(x),$
and we are done. ∎
Theorem 4.6.
Let $(X,f)$ be a compact dynamical system. Then there is a chain-recurrent point $x\in X$ such that $x\,\mathcal{C}_{\subseteq}\,x$.
Proof.
Pick $x\in X$. Let $\lambda$ be an ordinal such that $|\lambda|>|X|$ and $S_{\lambda}(x)$ a set defined by means of the transfinite procedure described in Step 1 of the proof of Theorem 3.1. Let $\alpha\leq\lambda$ be the least ordinal greater than 0 such that $x_{\alpha}\in S_{\alpha}(x)$, which means that $x_{\alpha}=x_{\beta}=:z$ for some $0\leq\beta<\alpha$. Therefore, we have $z\in CR_{f}$. Applying Lemma 4.5 with $\eta=\beta$ and $\xi=\alpha$, we obtain that $z\,\mathcal{C}_{\subseteq}\,z$. ∎
Definition 4.7.
Let $(X,f)$ be a dynamical system, and $M\subseteq X$. We say that $M$ is $\mathcal{C}_{\subseteq}$-transitive if for every $x,y\in M$ we have that $x\,\mathcal{C}_{\subseteq}\,y$.
We now prove that every compact system has a $\mathcal{C}_{\subseteq}$-transitive subsystem. For this, we need a preliminary lemma.
Lemma 4.8.
Let $(X,f)$ be a compact dynamical system and $M$ be a subset of $X$. If $M$ is internally chain transitive, then $M$ is $\mathcal{C}_{\subseteq}$-transitive.
Proof.
Let $d$ be a metric compatible with the topology on $X$. Pick $x,y\in M$ and let $\{\epsilon_{m}\}_{m\in\mathbb{N}}$ be a non-increasing sequence of positive reals converging to zero. Since $M$ is internally chain transitive, there is an $\epsilon_{1}$-chain $C_{\epsilon_{1}}\subseteq M$ from $x$ to $y$.
Proceeding by induction, assume that, for every positive integer $m\leq n$, there is an $\epsilon_{m}$-chain $C_{\epsilon_{m}}\subseteq M$ from $x$ to $y$, and that, for all $m<n$, one has $C_{\epsilon_{m}}\subseteq C_{\epsilon_{m+1}}$. Assume that $C_{\epsilon_{n}}$ consists of the points:
$x=x_{0},x_{1},\dots,x_{k}=y.$
Since $C_{\epsilon_{n}}\subseteq M$ and $M$ is internally chain transitive, for every $0\leq j<k$, there is an $\epsilon_{n+1}$-chain $C_{\epsilon_{n+1}}^{j}\subseteq M$ from $x_{j}$ to $x_{j+1}$, consisting of the points:
$x_{j}=x_{j_{1}},x_{j_{2}},\dots,x_{j_{h_{j}}}=x_{j+1}.$
Therefore, the points
$x=x_{0}=x_{{0}_{1}},x_{{0}_{2}},\dots,x_{{0}_{h_{0}}}=x_{{1}_{1}},\dots,x_{{1}_{h_{1}}}=x_{{2}_{1}},\dots\dots\dots,x_{{k-1}_{h_{k-1}}}=x_{k}=y$
form an $\epsilon_{n+1}$-chain $C_{\epsilon_{n+1}}$ from $x$ to $y$ such that
$C_{\epsilon_{n}}\subseteq C_{\epsilon_{n+1}}=\bigcup_{j=0}^{k-1}C_{\epsilon_{n+1}}^{j}\subseteq M,$
and we are done. ∎
In general, the converse implication does not hold, as elementary examples show. Consider indeed the set $S=\{0,1,\ldots,n\}$ with $n\in\mathbb{N}$ and the map $f:S\rightarrow S$ defined as $f(x)=(x+1)\mod n$. Then, for any proper subset $M\subsetneq S$ we see that $M$ is $\mathcal{C}_{\subseteq}$-transitive but not internally chain transitive. It is also easy to give examples in which $M$ is invariant. For instance, consider the logistic map with parameter 4 (that is the system $([0,1],f$, with $f(x)=4x(1-x))$) and take $M=\{0,\frac{3}{4}\}$, or the identity map on $[0,1]$ and take $M$ equal to any subset of $[0,1]$ containing an isolated point in the relative topology. Since $\mathcal{C}_{\subseteq}\subseteq\mathcal{C}$, this also shows that in general chain-transitivity does not imply internal chain transitivity.
Corollary 4.9.
Let $(X,f)$ be a compact dynamical system. If $X$ is $\mathcal{C}$-transitive, then $\mathcal{C}_{\subseteq}=\mathcal{C}$.
Proof.
Since $\mathcal{C}=X^{2}$, we have that $X$ is internally chain transitive. By Lemma 4.8, $X$ is $\mathcal{C}_{\subseteq}$-transitive, which means that $\mathcal{C}_{\subseteq}=X^{2}=\mathcal{C}$. ∎
Theorem 4.10.
Let $X$ be a compact dynamical system. Then $(X,f)$ has a $\mathcal{C}_{\subseteq}$-transitive subsystem.
Proof.
We will use the notation defined in the proof of Theorem 3.9.
By Theorem 3.9, the dynamical system $(X,f)$ has a $\mathcal{C}$-transitive subsystem. More precisely, there exists a least ordinal $\nu$ such that the set $\{x^{\gamma}\}_{\gamma<\nu}$ is defined, while it is impossible to define the point $x^{\nu}$.
Let us assume that $\nu$ is a successor ordinal. Then one of the following verifies:
  1. 1.
    The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is a closed, invariant and internally chain transitive subset. By Lemma 4.8, it follows that $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is $\mathcal{C}_{\subseteq}$-transitive and we are done.
  2. 2.
    The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is not closed and there exists an ordinal $\xi<\nu$ such that the set $M_{\xi,\nu}$ is closed, invariant and internally chain transitive. By Lemma 4.8, it follows that $M_{\xi,\nu}$ is $\mathcal{C}_{\subseteq}$-transitive and we conclude.
On the other hand, assume that $\nu$ is a limit ordinal. By the argument in the proof of Theorem 3.9, if we cannot define the point $x^{\nu}$, then there exists an ordinal $\xi<\nu$ such that $M_{\xi,\nu}$ is a closed, invariant and internally chain transitive subset, that is a chain transitive subsystem. By Lemma 4.8, it follows that $M_{\xi,\nu}$ is $\mathcal{C}_{\subseteq}$-transitive and we conclude. ∎
The relation $\mathcal{C}_{\subseteq}$ is clearly transitive. In the next example, we show that the relation $\mathcal{C}_{\subseteq}$ is, in general, not closed even if the map is continuous.
Example 4.11.
Let $\Sigma\subseteq\{0,1\}^{\omega}$ be the subshift defined by the following countable set of forbidden words:
$\{01^{j}0\text{ such that }j\geq 2\}\quad\cup\quad\{1^{j}0^{h}\text{ such that}j>h\}.$
Set $w_{k}:=1^{k}0^{k}10^{\infty}.$ The subshift $\Sigma$ is given by the words
$\{\sigma^{j}(w_{k}):j,k\in\mathbb{N}_{0}\}\cup\{1^{\infty}\}\cup\{0^{\infty}\}$
since
$\lim_{k\to\infty}w_{k}=1^{\infty}\quad,\quad\lim_{k\to\infty}\sigma^{k}w_{k}=0^{\infty}.$
[…]
$\mathcal{O}(x_{1})=\{1^{k-j}0^{p}10^{\infty}\}_{j=1}^{k-1}\cup\{0^{p-j}10^{\infty}\}_{j=0}^{p}\cup\{0^{\infty}\}.$
Since $p\geq k\geq 2m+1$, we have that $11^{h}0^{m}$ 1 is not a subword of any element of $\mathcal{O}(x_{1})$, thus we have:
$d(u,\mathcal{O}(x_{1}))\geq 2^{-2m}>\epsilon_{2m}.$
By the definition of $\Sigma$, it is clear that $(x,u)\notin\mathcal{C}_{\epsilon_{2m}}$ for all $x\in\mathcal{O}(x_{1})$. Therefore, $u\notin C_{\epsilon_{2m}}$ and we can conclude.
In the table below we summarize our main results (right) and their continuous versions (left).
Assuming $f$ continuous No assumptions on $f$
There exists an almost periodic point There exists a $\mathcal{C}$-recurrent point
(in fact also $\mathcal{C}_{\subseteq}$-recurrent and belonging to $GR(f)$)
Every point is in recurrence relation with a $\mathcal{C}$-recurrent point Every point is in $\mathcal{C}$-recurrence relation with a $\mathcal{C}$-recurrent point
There exists a minimal subsystem There exists a $\mathcal{C}$-transitive subsystem
Every point is in recurrence relation with a $\mathcal{C}$-transitive subsystem Every point is in $\mathcal{C}$-recurrence relation with a $\mathcal{C}$-transitive subsystem
Table 1: Comparison of existence results in compact systems with continuity and with no assumptions on the map.